# JANUS : A Difference-Oriented Analyzer For Financial Centralization Risks in Smart Contracts

### v1

@@ Abstract
Some smart contracts violate decentralization principles by defining privileged accounts that manage other users’ assets without permission, introducing centralization risks that have caused financial losses. Existing methods, however, face challenges in accurately detecting diverse centralization risks due to their dependence on predefined behavior patterns. In this paper, we propose JANUS, an automated analyzer for Solidity smart contracts that detects financial centralization risks independently of their specific behaviors. JANUS identifies differences between states reached by privileged and ordinary accounts, and analyzes whether these differences are finance-related. Focusing on the impact of risks rather than behaviors, JANUS achieves improved accuracy compared to existing tools and can uncover centralization risks with unknown patterns.
To evaluate JANUS’s performance, we compare it with other tools using a dataset of 540 contracts. Our evaluation demonstrates that JANUS outperforms representative tools in terms of detection accuracy for financial centralization risks . Additionally, we evaluate JANUS on a real-world dataset of 33,151 contracts, successfully identifying two types of risks that other tools fail to detect. We also prove that the state traversal method and variable summaries, which are used in JANUS to reduce the number of states to be compared, do not introduce false alarms or omissions in detection.
Index Terms:
Smart Contract, Symbolic Execution, Centralization Risk.
@@ I Introduction
Solidity, one of the most popular smart contract languages [1], is supported by several blockchain platforms, e.g., Ethereum [2] and Binance Smart Chain [3]. Solidity smart contracts have been widely adopted, managing digital assets valued at over \$10 billion [4]. However, some smart contracts introduce the risk of centralization in their attempt to facilitate the management of users (or claim to do so). For instance, the SoarCoin contract included a special function allowing the token issuer to retrieve tokens at no cost. Although developers claimed this feature was intended for airdrops and development activities, its misuse in June 2018 resulted in an Australian company losing assets worth \$6.6 million [5]. A similar incident occurred with the Squid token in November 2021, causing financial losses of \$3.38 million for contract users [6]. These cases highlight a common issue: contracts defining privileged accounts that can manage other users’ assets without permission, thereby introducing centralization risks. Such risks not only pose potential threats to contract users but can also lead to financial losses. Consequently, there is an urgent need for methods to detect and mitigate contract centralization risks to protect users’ assets.
Researchers have proposed automated detection methods for centralization risks in smart contracts. Pied-piper [7] identifies five common patterns of code associated with centralization risks. It employs static Datalog analysis [8] to detect contracts conforming to these patterns, followed by directed fuzzy testing [9] to minimize false positives. Similarly, Tokeer [10] focuses on transfer-related functional modules in contract code. It generates oracles based on four known rug pull contract patterns and uses Datalog analysis for detection. Rug pull contracts [11], a type of malicious fraud contract, often incorporate code for asset transfer with centralization risks. Consequently, detection methods developed for rug pull contracts can also be applied to identify centralization risks in smart contracts.
However, existing methods face challenges in accurately detecting diverse code with centralization risks in real-world scenarios due to their reliance on predefined behavior patterns. They may underreport when faced with unknown patterns or variants of known behavior patterns. Additionally, inaccurate predefined patterns can lead to misclassification of secure contracts as risky ones.
@@
In this paper, we propose a detection framework for centralization risks in smart contracts independently of specific behavior patterns, offering improved accuracy compared to existing methods and the ability to uncover risky codes of unknown patterns.
First, our observation reveals that existing risky code often leverages privileged accounts to manipulate variables through execution paths inaccessible to ordinary accounts. Privileged accounts are those designated by developers for management purposes, such as the contract owner. This paper aims to detect centralization risks associated with these privileged accounts to avoid financial losses for users Therefore, we focus on manipulation made by these accounts to financial variables, e.g., tokens or ether balances, related to users’ assets.
Based on this observation, we propose an automated detection framework for centralization risks, as illustrated in Fig. 1. The framework obtains differences between states reached by privileged and ordinary accounts through an iterative algorithm (shown in the dotted box) and analyzes whether these differences relate to financial variables. If so, it indicates that privileged accounts can perform specialized operations on their own or ordinary accounts’ financial assets, suggesting the possible presence of risks. By focusing on the ultimate impact of risks rather than specific behavior patterns, our framework enables more accurate detection than existing tools.
@@
    Given a contract, symbolically execute the functions within it using both privileged and ordinary accounts, then collect the differences in outcome states into an initial difference set.
@@
Finally, we implement an automated detection tool JANUS [13] for financial centralization risks based on our framework and evaluate its effectiveness. We collect a comparison dataset containing 540 smart contracts, and the experimental results on this dataset show that JANUS is more accurate than existing tools, Pied-piper and Tokeer. In addition, we collect 33,151 real-world smart contracts and JANUS identifies 8,391 smart contracts with potential risks, finding two kinds of risks that cannot be detected by existing tools.
@@
  • We propose a framework for detecting financial centralization risks based on the differences in contract states reached by privileged and ordinary accounts, which does not depend on the specific behavior patterns of risky codes.
  • We propose an iterative algorithm for obtaining function result differences, using a difference-oriented state traversal method to optimize contract state space exploration and using variable summaries to ensure algorithm termination. We provide proofs that these methods do not introduce false alarms or cause omissions in risk detection.
@@
  • We also collect 33,151 real-world smart contracts and use JANUS to find 8,391 smart contracts with centralization risks, and find two kinds of risks with unknown patterns that cannot be detected by existing tools.
@@ II-A Solidity Smart Contracts
II-B Existing Centralization Risks in Contracts
The design principle of blockchain and smart contracts is decentralization, which aims to reduce the number of entities that users need to trust [15]. However, some existing contracts define privileged accounts, reintroducing centralized management. This leads to the risk of exploitation by malicious developers who could attack users, as there is no guarantee of the contract developers’ trustworthiness. In this paper, we refer to these risks as centralization risks and classify them into five categories based on their potential consequences [7]:
@@ III-A Overview
Based on these features, we propose the framework shown in Fig. 1 to detect centralization risks:
@@ III-B Technical Challenges
First, our framework faces the state space explosion problem. As illustrated in Fig. LABEL:fig:states, each function execution using privileged accounts and ordinary accounts, denoted by arrows with subscripts p and o respectively, results in two states. To reach any two states in the same layer of the state space separately, at least one function must be executed using accounts with different permissions. Consequently, to fully capture the impact of different accounts on the states, we need to compare the states in each layer of the state space pairwise. For example, in Layer 4 of Fig. LABEL:fig:states, 28 comparisons (denoted by dotted lines) are required for 8 states. As the number of functions increases, the total number of states grows exponentially, along with the number of comparisons. This exponential growth may ultimately prevent our method from completing state traversal and comparison within limited resources.
@@
Output: differences bewtween the states reached by executing $C$ with privileged account $a_{p}$ and ordinary account $a_{o}$
1:$F\leftarrow$ the functions in $C$
2:$s_{0}\leftarrow$ initial state after executing the constructor of $C$
3:$D\leftarrow\emptyset$
4:$S_{next}\leftarrow\emptyset$
@@ III-C Our Iterative Algorithm
To address the aforementioned challenges, we propose two methods used in the iterative algorithm: difference-oriented state traversal and variable summaries. Difference-oriented state traversal reduces the number of states to be traversed by focusing on a subset of state spaces. Variable summaries make the number of possible state differences finite by abstracting the information in the states.
@@
Algorithm 1 illustrates the overall flow of the iterative algorithm, which takes the smart contract source code as input and produces the differences between the states reached by privileged accounts and ordinary accounts as output. Note that the steps about state update, e.g. Line 11, will be introduced in part 2). The detailed flow is as follows:
a) Initialization. Retrieve all functions from the smart contract source code for analysis (Line 1). Obtain the initial state, $s_{0}$, by symbolically executing the contract’s constructor using a symbolic execution engine implemented based on Slither[16] and Z3 [17] (Line 2). Initialize an empty ordered set $D$ to store differences (Line 3).
b) Difference Set Initialization. For each function (Line 5), conduct symbolic execution using both privileged and ordinary accounts (Lines 6-7). Summarize and compare execution results to derive differences (Line 8). $\Delta$ outputs the difference between two states, it will be introduced in the parts about variable summary along with $\Phi$. Aggregate resultant differences from all functions into a set (Lines 9-10) for use in subsequent difference propagation. Here, function $add$ is used to add an element into a set.
c) Difference Propagation. Begin an iterative propagation process, continuing until no new differences are generated in the previous round (i.e., when $D_{new}$ is empty). For each difference in $D_{new}$, first search for related functions (Line 21). The difference between two states consists of the difference between the individual variables in the two states, and its related functions are the functions that depend on the variables whose values differ between the two states. We implement the search for related functions of differences using Slither. Then for each related function (Line 24), conduct symbolic execution using privileged and ordinary accounts (Lines 25-26) and compare execution results to obtain differences (Line 27). If there are newly discovered differences (Lines 33), add them to the $D_{new}$ set (Lines 34 and 39) for processing in subsequent rounds. Here, function $extend$ is used to combine two sets. Finally, this iterative process ends up with a converged set of differences.
@@
We categorize the differences between two states at any layer of the state space into two types: those caused solely by differences in account permissions, and those stemming from differences in preceding states. Leveraging these distinct categories, we introduce a difference-oriented state traversal method comprising two primary steps to reduce the number of states requiring traversal and comparison.
First, for each layer in the state space, we selectively compare two states resulting from the execution of privileged and ordinary accounts, respectively. As illustrated in Fig. LABEL:fig:states, we traverse and compare only the states represented by green nodes, thereby preserving the first category of differences. Specifically, we implement this step as follows:
a) In Difference Set Initialization of Algorithm 1, we collect only the state $s_{p}$ (resulting from privileged account executions) and add them to the $S_{next}$ set (line 11).
@@
b) In Difference Propagation of Algorithm 1, for each difference $delta$ in $D_{new}$, we use the corresponding state $S_{next}[D_{new}.index(delta)]$ as the starting point for the symbolic execution of current round. Here, $D_{new}.index(delta)$ represent the index of $delta$ in set $D_{new}$. For each function $f$ to be executed (Line 24), we also collect only the state $s_{p}$, adding it to both the temporary state set $S_{tmp}$ and $S_{next}$ as the initial state for the next round.
@@
Second, motivated by the taint analysis technique [18], we add labels for variables exhibiting differences during comparison. This method preserves the second category of differences without increasing the number of traversed nodes. For example, in Fig. LABEL:fig:states, when comparing states $s_{10}$ and $s_{11}$, we add labels to variables with different values. Subsequently, we conduct a data flow analysis of function $f_{2}$ and add labels for variables dependent on the labeled ones. If a labeled variable undergoes reassignment in $f_{2}$, and the assignment statement is independent of any labeled variable, we remove the label from that variable. In the comparison of states $s_{20}$ and $s_{21}$ in Layer 3, labeled variables are not directly compared but are instead incorporated into the set of difference variables. These labeled variables, which display value differences between states $s_{22}$ or $s_{23}$ and $s_{20}$, are included in the subsequent analysis of financial variables. To implement this step, we replace the states in Algorithm 1 with labeled state $l=(\sigma,\pi,\theta)$, where $\sigma$ is a map from variable names to values, $\pi$ denotes the sequence of statements executed since the last state, and $\theta$ is a map from variable names to boolean values indicating whether the variable is labeled. Similar to the transition of states, the transition from labeled state $l$ to $l^{\prime}$ is written as $l\stackrel{{\scriptstyle f(M)}}{{\longrightarrow}}_{x}l^{\prime}$. Due to the page limit, we present the modified algorithm in the Supplemental Material.
@@
Specifically, we summarize the information of a variable into a set of key-value pairs, where different types of variables correspond to different numbers of keys. As shown in TABLE V, we design corresponding variable summaries for the following variable types:
@@
[…]
    Here, balances[_from] and balances[_to] are treated as Numeric Type variables, and their summaries are generated according to TABLE V. Function $\phi(l,l^{\prime},v)$ outputs the summary of variable $v$ under the transition from $l$ to $l^{\prime}$.
  • Other Types. Two additional types of variables, namely exec_state and ether, are automatically generated during the symbolic execution of contracts. The exec_state is a string-type variable that can take on three possible values: $success$, $revert$, or $selfdestruct$. These values respectively signify whether the function execution succeeds, reverts, or causes the contract to self-destruct. Each possible value of exec_state corresponds to a key in its summary. When exec_state takes on a specific value, the corresponding key in the summary is set to true, while the others remain false. During the executions, a variable ether of type mapping(address=>uint) is employed to record the ether balances of accounts. This variable is handled in the same way as the mapping variable described above.
@@
Furthermore, we categorize the key-value pairs in the summaries into two types: those generated based on variable values, and those generated based on the executed statements. For example, is_increased and is_decreased of Numeric Type variables fall into the former category, as they are determined by changes in the variable’s value. Conversely, related_const_var belongs to the latter category, as it needs to be generated according to data dependencies obtained from executed statements. The specific categorization of each key-value pair is shown in TABLE V.

### v2

# JANUS : A Difference-Oriented Analyzer For Financial Centralized Risks in Smart Contracts

@@ Abstract
Some smart contracts violate decentralization principles by defining privileged accounts that manage other users’ assets without permission, introducing centralized risks that have caused financial losses. Existing methods, however, face challenges in accurately detecting diverse centralized risks due to their dependence on predefined behavior patterns. In this paper, we propose JANUS, an automated analyzer for Solidity smart contracts that detects financial centralized risks independently of their specific behaviors. JANUS identifies differences between states reached by privileged and ordinary accounts, and analyzes whether these differences are finance-related. Focusing on the impact of risks rather than behaviors, JANUS achieves improved accuracy compared to existing tools and can uncover centralized risks with unknown patterns.
To evaluate JANUS’s performance, we compare it with other tools using a dataset of 540 contracts. Our evaluation demonstrates that JANUS outperforms representative tools in terms of detection accuracy for financial centralized risks. Additionally, we evaluate JANUS on a real-world dataset of 33,151 contracts, successfully identifying two types of risks that other tools fail to detect. We also prove that the state traversal method and variable summaries, which are used in JANUS to reduce the number of states to be compared, do not introduce false alarms or omissions in detection.
@@ I Introduction
Solidity, one of the most popular smart contract languages [21], is supported by several blockchain platforms, e.g., Ethereum [16] and Binance Smart Chain [15]. Solidity smart contracts have been widely adopted, managing digital assets valued at over \$10 billion [41]. However, some smart contracts introduce the risk of centralization in their attempt to facilitate the management of users (or claim to do so). For instance, the SoarCoin contract included a special function allowing the token issuer to retrieve tokens at no cost. Although developers claimed this feature was intended for airdrops and development activities, its misuse in June 2018 resulted in an Australian company losing assets worth \$6.6 million [5]. A similar incident occurred with the Squid token in November 2021, causing financial losses of \$3.38 million for contract users [30]. These cases highlight a common issue: contracts defining privileged accounts that can manage other users’ assets without permission, thereby introducing centralized risks. Such risks not only pose potential threats to contract users but can also lead to financial losses. Consequently, there is an urgent need for methods to detect and mitigate contract centralized risks to protect users’ assets.
Researchers have proposed automated detection methods for centralized risks in smart contracts. Pied-piper [27] identifies five common patterns of code associated with centralized risks. It employs static Datalog analysis [22] to detect contracts conforming to these patterns, followed by directed fuzzy testing [6] to minimize false positives. Similarly, Tokeer [54] focuses on transfer-related functional modules in contract code. It generates oracles based on known rug pull contract patterns and uses Datalog analysis for detection. Rug pull contracts [37], a type of fraud contracts, often incorporate code for asset transfer with centralized risks. Consequently, detection methods developed for rug pull contracts can also be applied to identify centralized risks in smart contracts.
However, existing methods face challenges in accurately detecting diverse code with centralized risks in real-world scenarios due to their reliance on predefined behavior patterns. They may underreport when faced with unknown patterns or variants of known behavior patterns. Additionally, inaccurate predefined patterns can lead to misclassification of secure contracts as risky ones.
@@
In this paper, we propose a detection framework for centralized risks in smart contracts independently of specific behavior patterns, offering improved accuracy compared to existing methods and the ability to uncover risky codes of unknown patterns.
First, our observation reveals that existing risky code often leverages privileged accounts to manipulate variables through execution paths inaccessible to ordinary accounts. Privileged accounts are those designated by developers for management purposes, such as the contract owner. This paper aims to detect centralized risks associated with these privileged accounts to avoid financial losses for users. Therefore, we focus on manipulation made by these accounts to financial variables used to represent digital asset balances, e.g., the commonly used variable balanceOf indicating token balances.
Based on this observation, we propose an automated detection framework for centralized risks, as illustrated in Fig. 1. The framework obtains differences between states reached by privileged and ordinary accounts through an iterative algorithm (shown in the dotted box) and analyzes whether these differences relate to financial variables. If so, it indicates that privileged accounts can perform specialized operations on their own or ordinary accounts’ financial assets, suggesting the possible presence of risks. By focusing on the ultimate impact of risks rather than specific behavior patterns, our framework enables more accurate detection than existing tools.
@@
    Given a contract, symbolically execute the functions within it using both privileged and ordinary accounts, then collect the differences in outcomes into an initial difference set.
@@
Finally, we implement an automated detection tool JANUS [35] for financial centralized risks based on our framework and evaluate its effectiveness. We collect a comparison dataset containing 540 smart contracts, and the experimental results on this dataset show that JANUS is more accurate than existing tools, Pied-piper and Tokeer. In addition, we evaluate JANUS in 33,151 real-world contracts, finding two kinds of risks that cannot be detected by existing tools.
@@
  • We propose a framework for detecting financial centralized risks based on the differences in contract states reached by privileged and ordinary accounts, which does not depend on the specific behavior patterns of risks.
  • We propose an iterative algorithm for obtaining function result differences, using a difference-oriented state traversal method to optimize contract state space exploration and using variable summaries to ensure algorithm termination. We provide proofs that these methods do not introduce false alarms or omissions in risk detection.
@@
  • We evaluate JANUS on 33,151 real-world smart contracts. Manual inspection of 1,000 contracts confirms that it achieves over 90% precision and recall, while identifying two previously undetectable risk patterns missed by existing tools.
@@ II-A Solidity Smart Contracts
II-B Existing Centralized Risks in Contracts
The centralized risks, as presented in [27] and [48], are defined as functions that can be exclusively invoked by privileged accounts that may affect other accounts without notification or authorization. Based on summaries of security incidents and feedback from smart contract developers, these papers identify various centralized risks. Our paper focuses on five such risks that can cause financial losses to contract users:
@@ II-B Existing Centralized Risks in Contracts
Note that while [27] characterizes these risks as backdoor threats, we adopt the term ‘centralized risk’ proposed in [48] throughout our paper.
@@ III-A Overview
Based on these features, we propose the framework shown in Fig. 1 to detect centralized risks:
@@ III-B Technical Challenges
First, our framework faces the state space explosion problem. As illustrated in Fig. 8, each function execution using privileged accounts and ordinary accounts, denoted by arrows with subscripts p and o respectively, results in two states. To reach any two states in the same layer of the state space separately, at least one function must be executed using accounts with different permissions. Consequently, to fully capture the impact of different accounts on the states, we need to compare the states in each layer of the state space pairwise. For example, in Layer 4 of Fig. 8, 28 comparisons (denoted by dotted lines) are required for 8 states. As the number of functions increases, the number of states grows exponentially, along with the number of comparisons. This exponential growth may ultimately prevent our method from completing state traversal and comparison within limited resources.
Figure 8: An example of the state space explosion problem of our iterative algorithm.
@@
Output: differences between the states reached by executing $C$ with privileged account $a_{p}$ and ordinary account $a_{o}$
[…]
@@ III-C Our Iterative Algorithm
To address the aforementioned challenges, we propose two methods used in the iterative algorithm: difference-oriented state traversal and variable summaries. In the following, we first introduce the overall flow of the algorithm, then explain how difference-oriented state traversal is performed within this flow, and finally present the role and categorization of variable summaries used in the algorithm.
@@
Algorithm 1 illustrates the overall flow of the iterative algorithm, which takes the smart contract source code as input and produces the differences between the states reached by privileged accounts and ordinary accounts as output. The detailed flow is as follows:
  • Initialization. Retrieve all functions from the smart contract source code for analysis (Line 1). Obtain the initial state, $s_{0}$, by symbolically executing the contract’s constructor using a symbolic execution engine implemented based on Slither[13] and Z3 [50] (Line 2). Initialize an empty ordered set $D$ to store differences (Line 3).
  • Difference Set Initialization. For each function (Line 5), conduct symbolic execution using both privileged and ordinary accounts (Lines 6-7). Summarize and compare execution results to derive differences (Line 8). $\Delta$ outputs the difference between two states, it will be introduced in the parts about variable summary along with $\Phi$. Aggregate resultant differences from all functions into a set (Lines 9-10) for use in subsequent difference propagation. Here, function addadd is used to add an element into a set.
  • Difference Propagation. Begin an iterative propagation process, continuing until no new differences are generated in the previous round (i.e., when $D_{new}$ is empty). For each difference in $D_{new}$, first search for related functions (Line 21). The difference between two states consists of the difference between the individual variables in the two states, and its related functions are the functions that depend on the variables whose values differ between the two states. We implement the search for related functions of differences using Slither. Then for each related function (Line 24), conduct symbolic execution using privileged and ordinary accounts (Lines 25-26) and compare execution results to obtain differences (Line 27). If there are newly discovered differences (Lines 33), add them to the $D_{new}$ set (Lines 34 and 39) for processing in subsequent rounds. Here, extendextend is used to combine two sets. Finally, this iteration ends up with a converged difference set.
@@
We categorize the differences between two states at any layer of the state space into two types: those caused solely by differences in account permissions, and those stemming from differences in preceding states. Leveraging these distinct categories, we introduce a difference-oriented state traversal method to reduce the number of states requiring traversal by combining pruning and difference labeling while ensuring no loss of essential differential information.
@@
First, for each layer in the state space, we selectively compare two states resulting from the execution of privileged and ordinary accounts, respectively. Specifically, we implement this step as follows: In Difference Set Initialization of Algorithm 1, we collect only the state $s_{p}$ (resulting from privileged account executions) and add them to the $S_{next}$ set (line 11). In Difference Propagation of Algorithm 1, for each difference deltadelta in $D_{new}$, we use the corresponding state $S_{next}[D_{new}.index(delta)]$ as the starting point for the symbolic execution of current round. Here, $D_{new}.index(delta)$ represent the index of deltadelta in set $D_{new}$. For each function $f$ to be executed (Line 24), we also collect only the state $s_{p}$, adding it to both the temporary state set $S_{tmp}$ and $S_{next}$ as the initial state for the next round. Differences in subsequent states of $s_{o}$ are preserved through labeled variables.
@@
Second, motivated by the taint analysis technique [28], we add labels for variables exhibiting differences during comparison. This method preserves the second category of differences without increasing the number of traversed nodes. To implement this step, we replace the states in Algorithm 1 with labeled state $l=(\sigma,\pi,\theta)$, where $\sigma$ is a map from variable names to values, $\pi$ denotes the sequence of statements executed since the last state, and $\theta$ is a map from variable names to boolean values indicating whether the variable is labeled. Similar to the transition of states, the transition from labeled state $l$ to $l^{\prime}$ is written as $l\stackrel{{\scriptstyle f(M)}}{{\longrightarrow}}_{x}l^{\prime}$. Due to the page limit, we present the modified algorithm in [45]. Formal proofs for this mechanism are provided in Section III-E.
Taking Fig. 8 as an example, we traverse and compare only the states represented by green nodes, thereby preserving the first category of differences. Then, to preserve the second category of differences, when comparing states $s_{10}$ and $s_{11}$, we add labels to variables with different values. Subsequently, we conduct a data flow analysis of function $f_{2}$ and add labels for variables dependent on the labeled ones. If a labeled variable undergoes reassignment in $f_{2}$, and the assignment statement is independent of any labeled variable, we remove the label from that variable. In the comparison of states $s_{20}$ and $s_{21}$ in Layer 3, labeled variables are not directly compared but are instead incorporated into the set of difference variables. These labeled variables, which display value differences between states $s_{22}$ or $s_{23}$ and $s_{20}$, are included in the subsequent analysis of financial variables. Note that although $s_{22}$ and $s_{23}$ remain untraversed, labeling difference variables during the comparison of $s_{10}$ and $s_{11}$ ensures that when these nodes exhibit variable differences with other nodes in the same layer, the comparison of $s_{20}$ and $s_{21}$ will produce identical differences.
@@
Specifically, we summarize the information of a variable into a set of key-value pairs, where different types of variables correspond to different numbers of keys. As shown in TABLE IV, we design corresponding variable summaries for the following variable types:
@@
    $\phi(l,l^{\prime},\texttt{balances})=\{\texttt{\_to}:\phi(l,l^{\prime},\texttt{balances[\_to]})$
    $,\texttt{\_from}:\phi(l,l^{\prime},\texttt{balances[\_from]})\}$
    Here, balances[_from] and balances[_to] are treated as Numeric Type variables, and their summaries are generated according to TABLE IV. Function $\phi(l,l^{\prime},v)$ outputs the summary of variable $v$ under the transition from $l$ to $l^{\prime}$.
  • Other Types. Two additional types of variables, namely exec_state and ether, are automatically generated during the symbolic execution of contracts. The exec_state is a string-type variable that can take on three possible values: successsuccess, revertrevert, or selfdestructselfdestruct. These values respectively signify whether the function execution succeeds, reverts, or causes the contract to self-destruct. Each possible value of exec_state corresponds to a key in its summary. When exec_state takes on a specific value, the corresponding key in the summary is set to true, while the others remain false. During the executions, a variable ether of type mapping(address=>uint) is employed to record the ether balances of accounts. This variable is handled in the same way as the mapping variable described above.
@@
Furthermore, we categorize the key-value pairs in the summaries into two types: those generated based on variable values, and those generated based on the executed statements. For example, is_increased and is_decreased of Numeric Type variables fall into the former category, as they are determined by changes in the variable’s value. Conversely, related_const_var belongs to the latter category, as it needs to be generated according to data dependencies obtained from executed statements. The specific categorization of each key-value pair is shown in TABLE IV. Since both the summary types and the number of state variables in a contract are fixed, the maximum number of elements in the final difference set is bounded, and the iterative algorithm is theoretically guaranteed to terminate. Consequently, we do not set a maximum iteration count for the algorithm.